OpenStudy (anonymous):

Solve the following system of equations for x : 4x+y=11 and x+y=2

OpenStudy (anonymous):

I need helpp on this onee ...

OpenStudy (anonymous):

subtract second eqn from first to get 3x = 9 x = 9/3 x = 3 now substitute this in first eqn to get 4(3)+y=11 y = 11-12 y = -1 so solution set is (3, -1)